### v2

# TESSERACT Collaboration

Abstract
[…]
preprint: AIP/123-QED
I Motivation
[…]
At present, the field lacks an understanding of the fundamental processes setting the background environment and energy resolution of such detectors. For nearly the past decade, low threshold phonon detectors have observed orders of magnitude more events below several hundred eV than expected from background radioactivity or cosmic ray interactions Adari et al. (2022); Baxter et al. (2025), events often called the “low energy excess” (or LEE). These excess events couple both directly to a detector’s sensor films (“singles” LEE), and more problematically, to a detector’s phonon system (“shared” LEE) Anthony-Petersen et al. (2025); Angloher et al. (2024a). Similarly, our collaboration has observed excess noise in our phonon sensors, consistent with sub-threshold events in the phonon system (“correlated” or “phonon” noise) and our phonon sensors themselves (“uncorrelated” noise) Anthony-Petersen et al. (2025).
Superconducting qubits and similar devices have observed quasiparticle bursts in their superconducting films Mannila et al. (2022) including correlated bursts seemingly caused by phonon bursts in the substrateYelton et al. (2025). While potential connections to LEE-like phenomena have been proposedBaxter et al. (2025), the precise link between observations in qubits and low threshold particle detectors remains unclear.
A variety of explanations have been proposed for observations in low threshold phonon detectors, including the relaxation of thermal stress in the metal films used to construct phonon sensorsRomani (2024), or defects in the detector’s bulk crystalNordlund et al. (2024) or on its surface. This work probes the scaling of these backgrounds and noise terms with the detector volume (or equivalently mass).
II Experimental Setup
[…]
Figure 1: (Top left) Photograph of our 1 mm (left) and 4 mm thick (right) 1 cm2 silicon detectors. (Top right) Detail on mask design for our phonon sensors (QETs). For scale, the QET fins have a radius of approximately 140 $\mu$m. (Bottom) Sketch of backgrounds and shot noise sources we observe in our devices. Shared LEE backgrounds and correlated phonon shot noise appear at approximately 4x the rate in the 4 mm detector compared to the 1 mm detector.
To calibrate our detectors’ response, we illuminate them from above with pulses of 450 nm (2.755 eV) photons, diffused from the tip of a fiber which runs from the detector cavity to a room temperature laser. We trigger calibration events on a logic level signal that rises simultaneously with the firing of the laser, and analyze the response in each channel using optimum filtersAnthony-Petersen et al. (2025); Chang et al. (2025). Plotting the height of the resulting events in each channel (Fig. 2 left) , we observe the same features as in Ref. Anthony-Petersen et al. (2025): most photons are absorbed in the detector phonon system, and produce approximately equal responses in both channels, while occasionally photons are directly absorbed into the aluminum phonon collection fins, producing a large response only in one channel. Because multiple photons may strike the detector in one laser pulse, we also observe a superposition of these event types.
To combine the responses of the two channels of one detector, we use a “2x1” multi-channel optimum filter which simultaneously fits the response in both channels, scaling them together by one amplitude Anthony-Petersen et al. (2025); Chang et al. (2025). In our final dataset, we obtain a baseline phonon energy resolution of 258.5$\pm$0.4 meV for this combined channel response, improving on the resolution of our previously world leading detectorAnthony-Petersen et al. (2025); Chang et al. (2025) and previous work by other groupsMichielin (2024); Angloher et al. (2024b).
We periodically calibrate our detector, and find that the energy resolution of both detectors monotonically improves over the course of our run (Fig. 2 Bottom Right). In principle, this could be due to two factors: an improvement in the sensors phonon collection efficiency over time, or an improvement of noise in the detector. Monitoring our detectors’ phonon collection efficiencies over time, we find that they are approximately constant (see supplementary material section A), implying that our noise environment is improving with time.
Figure 2: 1 mm device calibration. (Left) Two dimensional histogram of calibration events. Most calibration photons hit the substrate, causing even responses in both channels (main diagonal events). Due to the energy resolution of our detector, individual photons appear as quantized quasi-Gaussian groups. Occasionally, photons hit a phonon sensor (QET), causing a large response in that channel (offset events). Around 1.1 eV of each 2.755 eV photon is absorbed in our sensors due to our phonon collection efficiency. (Top Right) Photon calibration spectra taken on days 2, 5, and 12. Note the improvement on baseline resolution over time. (Bottom Right) Measured phonon energy resolutions in the right, left, and combined channels over time.
III Measurement and Modeling of Detector Noise
To characterize our detectors’ noise over time, we take 2 hours of continuous data following each calibration dataset, and remove periods with abnormally small bias power (from periods with abnormally high environmental vibration or after large energy depositions). Additionally, to select periods without above threshold events that would contaminate and bias our noise, we find the largest magnitude pulse in the entire 200 ms period using an optimal filter, and select periods for which this largest fit magnitude pulse has a negative amplitude, i.e. the largest event in the trace is a statistical fluctuation rather than a real energy deposition. In total, 3.9% of the 200ms traces pass these very strict noise selection criteria and are used to estimate the noise Cross power Spectral Density (CSD, see supplementary material section C for further discussion of our cuts). We convert this CSD into the power domain by applying a three-pole responsivity modelLindeman et al. (2011); Irwin and Hilton (2005) for each channel (i.e. $\partial P/\partial I(f)$) that is estimated from IV and $\frac{\partial I}{\partial V}$ measurements as in Ref. Anthony-Petersen et al. (2025).
Similar to our observations in Ref. Anthony-Petersen et al. (2025), we observe that our noise is significantly in excess of modeled TES noise, both in the individual channel Power Spectral Densities (PSDs), and in noise which is correlated between the two channels of the detector (which should be statistically consistent with zero). The broad spectrum correlated noise is consistent with shot noise from many sub-threshold phonon events in the substrate, i.e. it has the same shape and energy partition (50% left, 50% right) as substrate-interacting calibration photons (Fig. 3 Top Right). The magnitude of this broad correlated noise appears to scale with the detector volume, i.e. the correlated noise is four times larger in the 4 mm detector than in the 1 mm detector (Fig. 3 Top Right). The correlated noise peaks in the 100Hz-1kHz range are harmonics of the vibrational resonant modes of the hanging wire bond-detector system.
After removing these correlated noise terms, we find that there is an additional quasi-flat uncorrelated noise term, in significant excess of the expected level of thermal fluctuation noise in the TESs, but in some tension with the flat shot noise model of Ref. Anthony-Petersen et al. (2025) (see supplementary material section D for further discussion).
Figure 3: (Left) Noise in the left channels of the 1 mm (top, green) and 4 mm (bottom, blue) detectors. Our noise is well in excess of our modeled TES noise (gray, dotted), and is composed of correlated phonon noise and uncorrelated noise as in Ref. Anthony-Petersen et al. (2025). The peaks around 150 Hz are due to vibration coupled noise. (Top Right) Correlated noise (i.e. off diagonal CSD element $S_{LR}$) in the 4 mm and 1 mm detectors. The correlated noise in the 4 mm detector is 4 times as large as the 1 mm detector. The amplitude of the 1 mm correlated noise is fit in the orange highlighted region (to avoid vibration coupled peaks), while the 4 mm noise is fit in the orange and yellow regions. (Bottom Right) Uncorrelated noise (i.e. $S_{LL}$, with the modeled correlated noise subtracted) in the left channels of the 1 mm and 4 mm detectors, consistent with the modeled TES Johnson noise at high frequencies and an excess noise term that is approximately flat and consistent between the two channels at low frequencies (see section D for further discussion).
Repeating this measurement of the correlated phonon noise amplitude, we are able to plot the correlated noise measured in the detector as a function of time (see Fig. 4). As with the phonon energy resolution, we find that the correlated phonon noise drops with time, causing the resolution improvement with time.
We additionally find that the bias power (i.e. the power applied to the TES to keep it in transition) of our TESs increases over time (see Fig. 4). We assume that this increase over time is essentially entirely due to a shifting parasitic power over time, as shifts in bias power due to a cooling dilution fridge base (mixing chamber, MC) temperature should be very small (see supplementary material section B). As with the correlated noise, the shift in the bias power over time is approximately four times larger in the 4 mm detector than in the 1 mm detector.
The observed monotonic decreases in both the excess noise in the TES and the parasitic power incident on the TES suggest that one source or a set of related sources might cause both phenomena. As the simplest benchmark model, quanta of energy $\varepsilon$ occurring at Poisson-distributed times with an average rate $R(t)\propto t^{-\kappa}$ would induce parasitic bias powers and excess noise terms
$P_{1,par}=\varepsilon R(t)=\varepsilon R_{0}t^{-\kappa}$ (1)
$P_{4,par}=\varepsilon 4R(t)=\varepsilon 4R_{0}t^{-\kappa}$ (2)
$S_{1,ex}=2\varepsilon^{2}R(t)=2\varepsilon^{2}R_{0}t^{-\kappa}=2\varepsilon P_{1,par}$ (3)
$S_{4,ex}=8\varepsilon^{2}R(t)=8\varepsilon^{2}R_{0}t^{-\kappa}=2\varepsilon P_{4,par}$ (4)
where we assume the rate of events $R(t)$ follows a power law, and the rate scales with the detector thickness. (An exponential trend is not a good fit to the data.)
Figure 4: Correlated noise level (Top Left) and Left + Right channel bias powers (Bottom Left) in the 1 mm (green) and 4 mm (blue) detectors as a function of time. (Right) Correlated noise level vs. Left + Right bias powers. Dashed lines show fit models, representing shot noise that scales in rate with detector volume with a shot noise quantum of $\varepsilon=0.68\pm 0.38$ meV.
We plot the excess correlated noise and bias power as a function of time and each other, and find good agreement with this model, which is fit simultaneously to the 1 mm and 4 mm detector data assuming a factor of 4 rate difference. We find $\varepsilon=0.68\pm 0.38$ meV, and $\kappa=0.635\pm 0.009$. Certainly, there is no reason to assume that the shot noise events all have an identical energy. As discussed in supplementary material section E, $\epsilon$ is readily generalizable to processes which have a distribution of energy depositions, where $\varepsilon=\frac{<E^{2}>}{<E>}$. So long as the distribution is not e.g. double peaked, $\varepsilon$ can still be thought of as a characteristic energy scale of the process.
Figure 5: (Top) The mass normalized shared background rates in the 1 mm and 4 mm detectors, with backgrounds measured in the SPICE 1%Anthony-Petersen et al. (2025); Chang et al. (2025), CPDv1Fink et al. (2021); Alkhatib et al. (2021), CPDv2 (previously unpublished), and CRESST 0.35g Si Angloher et al. (2023) detectors. Broadly, all six detectors seem to observe the same backrounds normalizing for mass. Note that CPD and CRESST detectors are one channel detectors which cannot reject singlesAnthony-Petersen et al. (2025), the increased background rates near threshold could be due to singles, as well as noise triggers or additional LEE backgrounds. (Top insert) Detail of the 1 mm, 4 mm, and SPICE 1% detector backgrounds. The orange and green bands show the energy ranges in which time dependence was measured (see bottom right). (Left Bottom) Background event energy partitioning between left and right channels, showing single and shared events. Note the energy reconstruction assumes a shared pulse shape in each channel. (Right bottom) Rate of shared events over time in the 1 mm detector in the shaded 2.5-5 eV and 5-10 eV bins in the center top figure. Dashed line shows the (weighted) average rate in the 5-10 eV range, and a power law fit in the 2.5-5 eV bin.
IV Above Threshold Shared Backgrounds
In addition to this phonon shot noise from below threshold events, we observe a high rate of above threshold background events (often called the “Low Energy Excess,” or LEEAdari et al. (2022); Baxter et al. (2025)). To investigate these LEE events, we recorded 2 or 12 hour periods of continuous background data following each calibration run, and triggered this dataset offline using an optimum filter optimized to search for events with a channel energy partition and pulse shape consistent with that of substrate absorbed calibration photons as in Ref. Chang et al. (2025). In Fig. 5 bottom left, a scatter plot shows the energy deposited in each channel without a constraint on the partition. Identical to observations in Ref. Anthony-Petersen et al. (2025), the LEE backgrounds split into 2 classes: “shared” events with a equal channel partition and a pulse shape identical to substrate absorbed photons, and “singles” events where nearly all energy is deposited in a single phonon sensor channel. Singles have a pulse shape with a fast rise, but a slow fall consistent with an energy deposition that saturates a subset of the athermal phonon sensors that are read out in parallel within a channel. Each event is then fit with a multi-channel optimum filter whose pulse shape and channel partition are fixed to the average of each LEE class (singles left, singles right, shared), and classified according to lowest $\Delta\chi^{2}$.
We see that the rate of above-threshold shared background events scales with the detector mass, as it does for below threshold phonon shot noise (the y-axis has been normalized by substrate mass). However, the rate of these above threshold shared events appears to be constant with respect to time at high energies. At lower energies, where statistics are better, the rate of shared LEE events does appear to decrease over time.
As in Ref. Anthony-Petersen et al. (2025), our singles appear to originate within the aluminum QET films, and we additionally observe that their rate decreases with time (see further discussion in supplementary materials section F). These observations give additional support to the proposal that the dislocation mediated relaxation of thermal stress in these aluminum QET films is responsible for these singles. Anthony-Petersen et al. (2025); Romani (2024)
V Discussion
Our observation of above threshold backgrounds and phonon shot noise which scales with the mass of our detector substrate leads us to conclude that these phonon bursts originate within the detector substrate itself. Through comparisons to detectors with different masses, surface-to-volume ratios, and holding methods,Anthony-Petersen et al. (2025); Chang et al. (2025); Alkhatib et al. (2021); Angloher et al. (2023) we can reject possible LEE scalings with the sidewall area, total surface area, or the holding method, either above (see Fig. 5) or below threshold (see supplementary materials section H). We also conclusively reject the hypothesis that our aluminum sensor films constitute our dominant source of phonon bursts,Romani (2024) but note that such a source could contribute at a subdominant level.
Such a mass scaling suggests that defects in the detector substrate are responsible for LEE and excess phonon noise observations, and is compatible with the model proposed in Ref. Nordlund et al. (2024). The strong rate vs. time since cooldown dependence we observe in our noise data is easily explained by assuming that the relevant excited states are thermally populated at 300K ( >> 1 meV), and then tunnel to lower energy states at mK temperatures. The observation of time dependence in above threshold LEE rates, both here and previouslyAngloher et al. (2023), is more difficult to explain given their eV energy scales are much higher than 300K $k_{B}$. A more complex model, perhaps involving the avalanche-like relaxation of many lower energy states, is needed to explain this time dependence in above threshold data.
The low energy phonon bursts we observe may also constitute the quasiparticle bursts and “poisoning” observed in superconducting qubits and similar devices. We can estimate the reduced quasiparticle density $x_{qp}=n_{qp}/n_{cp}=n_{qp}/(4\times 10^{6}\mu\mathrm{m}^{-3})$Martinis (2021) created by such phonon bursts by assuming that a constant power density $\rho_{p}$ of phonon events originating in the substrate is evenly divided over a $\sim$ 100 nm thick ground and qubit layer which covers essentially all of one face of a superconducting qubit chip. We assume that quasiparticle dynamics in the qubit and ground plane are dominated by recombination (rather than trapping), and that the system is in steady state, such that
$\frac{dx_{qp}}{dt}=0=-rx_{qp}^{2}+g$ (5)
[…]
$x_{qp}=\sqrt{\frac{g}{r}}=\sqrt{\frac{\rho_{p}}{\Delta_{Al}n_{cp}}\frac{t_{chip}}{t_{Al}}\frac{1}{r}}\approx 7\times 10^{-8}$ (6)
which is similar to the reduced quasiparticle density observed by well shielded silicon devices Mannila et al. (2022). Our observations also agree with the phonon origin Yelton et al. (2025) and reduction in rate over time Mannila et al. (2022) of such bursts.
VI Acknowledgments
This work was supported in part by DOE Grants DE-SC0019319, DE-SC0025523 and DOE Quantum Information Science Enabled Discovery (QuantISED) for High Energy Physics (KA2401032). This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE 1106400. This material is based upon work supported by the Department of Energy National Nuclear Security Administration through the Nuclear Science and Security Consortium under Award Number(s) DE-NA0003180 and/or DE-NA0000979. Work at Lawrence Berkeley National Laboratory was supported by the U.S. DOE, Office of High Energy Physics, under Contract No. DEAC02-05CH11231. Work at Argonne is supported by the U.S. DOE, Office of High Energy Physics, under Contract No. DE-AC02-06CH11357. W.G. and Y.Q. acknowledge the support by the National High Magnetic Field Laboratory at Florida State University, which is supported by the National Science Foundation Cooperative Agreement No. DMR-2128556 and the state of Florida.
The authors have no conflicts to disclose. The data that support the findings of this study are available from the corresponding author upon reasonable request.
Supplementary Materials Section A Phonon Collection Efficiency
To monitor the detectors’ phonon collection efficiency over time, we measure the responsivity $\partial P/\partial I$ of each sensor using IV sweeps and $\partial I/\partial V$ measurements at the operating point Irwin and Hilton (2005). We then calculate an average current-domain trace for events where a single photon is absorbed in the detector’s phonon system $i(t)$. Combining this with the responsivity, we can then calculate the phonon collection efficiency
[…]
where $E_{\gamma}$ is the energy of the photon hitting the detector and $\mathcal{F}$ and $\mathcal{F}^{-1}$ are Fourier and inverse Fourier transforms, respectively. We find that our phonon detectors’ phonon collection efficiencies remain roughly constant over time (See Fig. 6).
Figure 6: Phonon collection efficiencies for fit templates, measured from the photon calibrations, as a function of time. These efficiencies are roughly constant over time, meaning that the resolution improvements over time are due to improvements in the detector noise environment. 1 $\sigma$ error bars are shown.
Supplementary Materials Section B Effects of Mixing Chamber Cooling
In the steady state, we assume that two terms effect the bias power $P_{0}$ of our devices: the cooling of the TES through electron-phonon coupling, and parasitic power from any number of sources incident on our TESs.
$P_{0}=K_{ep}(T_{c}^{5}-T_{bath}^{5})-P_{par}$ (8)
Here, $K_{ep}$ gives the electron-phonon coupling in a certain TES operating at a temperature $T_{c}$ and connected to a bath of temperature $T_{bath}$.
We operate our mixing chamber (MC) at far below the $T_{c}$ of our TESs, such that even relatively large shifts in our mixing chamber and therefore bath temperature should have minimal impact on our TES cooling. Our detector housing is connected to our MC with several very high force gold plated copper to gold plated copper joints, and our detector is thermally connected to our housing with multiple gold wire bonds, which should very efficiently thermalize our detectors. Given our MC temperature drops from about 20 to 6 mK over the course of our run, we would expect a roughly 1% cooling power shift, assuming our detectors are at essentially the same temperature as our MC.
Our large ( 44\>44x) noise shift over time gives additional evidence that the mixing chamber temperature does not dominate our system’s dynamics. As an absolute upper limit, if our sensor’s bath temperature decreased from the sensor $T_{c}$ to a temperature of zero over the course of our datataking campaign, we would expect a maximum of a factor of 2 shift in noise due to shifting Thermal Fluctuation Noise (TFN) caused by this implausibly large shift in the bath temperature. Watkins (2022) Given we observe a noise shift of more than twice that magnitude, and as such an enormous temperature shift is highly implausible in any case, we can safely discount the possibility that shifting bath temperature causes our noise shift. (The characteristic phonon pole and correlated noise provide additional evidence against this model.)
We conclude that the monotonic increase in TES bias power over time is dominated by a decrease in $P_{par}$, and that the observed behavior of our noise and bias power is not dominated by shifts in the temperature of our fridge or detector.
Supplementary Materials Section C Data Quality and Single/Shared Discrimination Cuts
As in Ref. Anthony-Petersen et al. (2025), we apply quality cuts to all data series to ensure the detector is operating stably at a well characterized bias point. Additionally, we apply two special purpose cuts: a cut to ensure that traces with a large event in them are not used to construct noise estimates, and a cut which discriminated between shared and singles events in our background data.
Briefly, we cut on pre-pulse baseline, slope (pre-pulse baseline minus post-pulse baseline), and a low frequency $\chi^{2}$ statistic (only considering frequencies below 50 kHz), using the same general philosophy as in Ref. Anthony-Petersen et al. (2025). Unlike in Ref. Anthony-Petersen et al. (2025), we see significant vibration coupled noise around several hundred Hz (see Fig. 7), and use the pre-pulse baseline and slope cuts to remove events in which this noise dominates. Our low frequency $\chi^{2}$ cut primarily removes pileup events, and events which were triggered at the incorrect time.
Figure 7: The baseline cut in a representative 2 hour dataset on the 1 mm detector’s left channel. The approximately minute scale variations in the detector baseline are due to shifting vibration noise in the detector, which appears at roughly intervals and intensities in all datasets. The scattered points with very high baselines are events where a pulse occurs in the pre-pulse region (e.g. pileup). In the intervals with no points, we stop taking continuous data, and characterize the detector state by taking IV and dIdV measurements to characterize the TES state and photon calibration data to characterize the detector response.
For the noise analysis, we apply an additional cut intended to remove traces with large pulses that might dominate the noise if not removed. We do this by finding the largest magnitude (positive or negative) event in the each channel’s trace by using an unconstrained 1x1 OF, and only passing traces for which the largest event is negative in both channels, i.e. the largest event is consistent with a statistical fluctuation rather than a true event. In this sense, we select traces which are consistent with being dominated by pure noise.
Figure 8: Detail of the PSD of the 1 mm left and right channels with and without the PSD amplitude cut applied. Note that between the vibration induced peaks, the PSD amplitude cut does not change the broadband phonon shot noise (see text). Also note that the vibration-coupled noise causing the sharp peaks appears to couple more strongly to the right channel in some peaks.
Figure 9: Average current domain noise (arbitrary scale) in the Right channel of the 1 mm detector between 160 and 190 Hz as a function of time, showing events passing (pink) and failing (blue) the “PSD Amplitude” cut in this frequency range.
For the noise analysis, we also apply a more selective “PSD amplitude” cut to reduce vibration-induced peaks in the observed PSD and CSD which were not cut by our relatively loose baseline and slope cuts (see Fig. 8). These narrow resonance vibration-induced peaks (Q $\sim 10^{4}$) can be directly excited by mechanically thumping the fridge, and slowly fade away when the pulse tube is temporarily turned off (we cannot run for more than approximately 10 minutes in this configuration, and therefore take data with the pulse tube on). We also observe variation over time in the noise environment (see Fig. 9), presumably as the vibration source shifts into and out of resonance with the fridge/hanging detector mechanical system. We also observe that these peaks are power coupled: they disappear when the TES channel is superconducting or normal and therefore insensitive to changes in the incident power (but still sensitive to current or electronically coupled noise). Interestingly, on specific peaks, the right channel of the 1 mm detector appears to be significantly more strongly coupled to the noise source than the left channel. We interpret this to mean that the mechanism injecting power into our TESs does not always couple through the detector’s phonon system (which would equally couple to the two channels), and instead at least partially couples directly to individual channels. One possible explanation would be that the electrical wire bonds connecting to each channel vibrate and deposit energy directly into the TESs, with different damping constants for the wire bonds to each channel.
To perform a more selective cut of these high-vibration periods of time, we average the PSD of individual traces in a set relatively narrow ($\sim$30 Hz wide) frequency bins, and cut events for which the average in any bin exceeds a threshold. By varying the level of this threshold, we can more or less aggressively remove periods of high vibration noise. Because of the relatively low passage of this cut, we only apply it during the noise analysis. The findings in this paper do not depend on whether or not this PSD amplitude cut is applied, e.g. the level of measured phonon noise is the same with or without the PSD amplitude cut, as is the shift in the phonon noise over time.
For the background analysis, we discriminate singles from shared events by comparing the $\chi^{2}$ values for 2x1 OFs with different event typologies (as in Ref. Anthony-Petersen et al. (2025)). For example, we select singles in the right channel by requiring that a right singles 2x1 OF has a lower $\chi^{2}$ value than either the best fit left singles 2x1 OF or the best fit shared 2x1 OF.
Supplementary Materials Section D Uncorrelated Noise Systematics
As in Ref. Anthony-Petersen et al. (2025), we subtract the correlated noise from the total noise to find the uncorrelated noise in Fig. 3. We do this subtraction using four methods, finding consistent results:
  1. 1.
    We make the assumption that the uncorrelated noise in the left and right channels of the same detector is identical, and that correlated noise couples to the two channels with a constant $k(f)$ that can vary as a function of frequency, from e.g. the presence of absence of a vibration peak.
    $S_{ll}(f)=U(f)+C(f)k(f)$ (9)
    $S_{rr}(f)=U(f)+C(f)/k(f)$ (10)
    $S_{lr}(f)=C(f)$ (11)
    $j(f)=\frac{S_{ll}-S_{rr}}{S_{lr}}=k(f)-1/k(f)$ (12)
    $k(f)=\frac{1}{2}\Big{(}\sqrt{j(f)^{2}+4}-j(f)\Big{)}$ (13)
    $U(f)=S_{ll}(f)-S_{lr}(f)k(f)$ (14)
    Here, $U(f)$ and $C(f)$ are the uncorrelated and correlated noise terms, and $S$ is the CSD with on diagonal elements $S_{ll},S_{rr}$ and off diagonal elements $S_{lr}$.
  2. 2.
    We assume that for phonon coupled noise, both channels couple identically to the correlated noise, so
    $S_{ll}(f)=U_{l}(f)+C(f)$ (15)
    $S_{rr}(f)=U_{r}(f)+C(f)$ (16)
    $S_{lr}(f)=C(f)$ (17)
    $U_{l}(f)=S_{ll}(f)-S_{lr}(f)$ (18)
    We assume that the coupling to each channel is identical (which slightly deviates from the measured coupling of the two channels in the photon calibration).
  3. 3.
    We model the correlated phonon noise by fitting a three pole (two fall time, one rise time) exponential pulse model to the photon calibration data in the power domain, and by transforming this analytic pulse model to the frequency domain to get templates $t_{ll}(f)$, $t_{lr}(f)$, $t_{rr}(f)$.
    $S_{ll}(f)=U_{l}(f)+mt_{ll}(f)$ (19)
    $S_{rr}(f)=U_{r}(f)+mt_{rr}(f)$ (20)
    $S_{lr}(f)=mt_{lr}(f)$ (21)
    where $m$ is an arbitrary scale factor we measure by fitting $t_{lr}(f)$ to $S_{lr}(f)$.
  4. 4.
    Similarly to method 3, we model the correlated phonon noise by averaging photon calibration events, and by transforming this averaged data-driven template to the power and frequency domain to get templates $t_{ll}(f)$, $t_{lr}(f)$, $t_{rr}(f)$.
    $S_{ll}(f)=U_{l}(f)+mt_{ll}(f)$ (22)
    $S_{rr}(f)=U_{r}(f)+mt_{rr}(f)$ (23)
    $S_{lr}(f)=mt_{lr}(f)$ (24)
    where $m$ is an arbitrary scale factor we measure by fitting $t_{lr}(f)$ to $S_{lr}(f)$. Methods 3 and 4 differ in that 3 uses an analytic model of the pulse shape, whereas 4 uses a data-driven averaged calibration pulse as the correlated noise model.
See Fig. 10 for a graphical comparison of these methods.
Figure 10: Uncorrelated noise in the left channel of the 1 mm detector, reconstructed using four different methods (see text). In Fig. 3, we use Method 1. Methods 3 and 4 are on top of each other at lower frequencies, producing a brown appearing line.
Figure 11: Uncorrelated noise in each channel, measured using method 4 (see above). Grey dashed line shows modeled TES noise for the 1 mm left channel, all other channels have similar modeled noise spectra. Black dashed lines show TES noise + an arbitrary broadband constant, showing that no channel seems compatible with this simple model proposed in Ref. Anthony-Petersen et al. (2025).
The uncorrelated noise obviously deviates from the TES noise + flat shot noise model we proposed in Ref. Anthony-Petersen et al. (2025), and deviates in seemingly opposite ways in the 1 mm and 4 mm detectors. Several possibilities might explain this deviation from the expected form:
  • Phonon collection position dependence that changes solely the relative partition of energy between the channels will produce uncorrelated shot noise with an average phonon pulse shape from subthreshold phonon LEE relaxation events. Optical calibration energy distributions are consistent with a $\sim$ 4% fractional energy position dependence in the 4 mm device. However, this likely overestimates the true position dependence for uniformly distributed events throughout the bulk of the substrate since the optical calibration photons are preferentially absorbed on the phonon sensor instrumented surface. In the 4 mm device, the excess measured uncorrelated noise is consistent with a $8\times 10^{-37}$ W2/Hz flat noise possibly from localized relaxation in the aluminum fins plus additional shot noise with a phonon shape due to a 2% fractional energy position dependence.
    The excess shot noise above the phonon collection bandwidth but below the TES sensor pole seen in the 1 mm device can not be understood with such a simple model. Potentially, models where the position dependence changes both the pulse shape and the energy partition between the 2 channels could model the rise in excess uncorrelated noise found at higher frequencies in the 1 mm device. This explanation is not entirely self-consistent; we both expect and see more position dependence in the 1 mm thick device in the photon calibration, yet see less uncorrelated noise at all frequencies.
  • Our power-to-current responsivity model ($dP/dI$) might be inaccurate. We use standard models for this responsivity Lindeman et al. (2011); Irwin and Hilton (2005) which accurately predict current-domain pulse shapes for energy impulses in our TESs from photons Romani et al. (2024), giving us some reassurance that we are accurately modeling TES responsivity. However, our QETs (which have a different structure from TES photon sensors) could have additional thermal dynamics that do not couple as expected to our $\partial I/\partial V$ excitations. For example, “third” poles seen in our $\partial I/\partial V$ could couple to current through the TES in addition to the thermal couplings proposed in Ref. Lindeman et al. (2011). Directly probing our TESs’ $dP/dI$ through e.g. exciting our TESs with a low energy ($\sim 300$ meV) photon source could probe this potential systematic.
  • We do not attempt to model the Internal Thermal Fluctuation Noise (ITFN) Lindeman et al. (2011) caused by stocastic energy flucations between the independent thermal degrees of freedom in our 3 pole TES dynamical model. Given that our observed noise (even after “decorrelation”) dominated our modeled TFN noise by close to an order of magnitude, we find it unlikely that this ITFN noise due to additional heat capacities could be dominant.
Ultimately, we cannot pinpoint the origin of this effect, and are therefore reluctant to draw strong conclusions about the origin of our uncorrelated noise. We do note that the uncorrelated noise is roughly equal in magnitude for both 1 mm and 4 mm channels around 3-5 kHz, where neither correlated phonon noise nor TES noise should be dominant. This suggests that a common noise source (e.g. originating within the sensors themselves) causes this noise in all channels, however, without an accurate noise model, we feel that this observation is inconclusive.
Without a well supported model for this uncorrelated noise, we fear that attempts to measure changes in the uncorrelated noise over time will be subject to unknown systematics. Attempting to do this measurement by subtracting the TES noise and averaging the uncorrelated noise between 3 and 5 kHz (where presumably there is not too much interference from either phonon or TES noise) does appear to show a reduction in the uncorrelated noise over time (see Fig. 14). However, we are unable to draw strong conclusions about this apparent effect, and caution that it is likely due to systematic effects (e.g. phonon noise which we cannot subtract due to position dependence) rather than a true reduction over time in the uncorrelated noise.
Supplementary Materials Section E Shot Noise From Energy Spectra
Generically, we can extend the results of Eqns. 1 and 3 to account for arbitrary event spectra $\frac{\partial R}{\partial E}(E)$ by integrating
[…]
where $\varepsilon_{l}$ is the lowest energy event the detector is sensitive to (set by e.g. the superconducting bandgap $\Delta_{Al}$).
For an exponential event distribution
$\frac{\partial R}{\partial E}(E)=\frac{R_{0}}{\varepsilon_{0}}e^{-E/\varepsilon_{0}}$ (27)
we find
$P_{par}=R_{0}e^{-\varepsilon_{l}/\varepsilon_{0}}(\varepsilon_{l}+\varepsilon_{0})$ (28)
$S_{p}=2R_{0}e^{-\varepsilon_{l}/\varepsilon_{0}}(2\varepsilon_{0}^{2}+2\varepsilon_{0}\varepsilon_{l}+\varepsilon_{l}^{2})$ (29)
$\frac{S_{p}}{2P_{par}}=\frac{2\varepsilon_{0}^{2}+2\varepsilon_{0}\varepsilon_{l}+\varepsilon_{l}^{2}}{(\varepsilon_{l}+\varepsilon_{0})}$ (30)
Fitting a line to the $S_{p}$ vs. $P_{par}$ data, we would fit $2\varepsilon_{0}$ or $\varepsilon_{l}$ for $\varepsilon_{l}>>\varepsilon_{0}$ and $\varepsilon_{l}<<\varepsilon_{0}$ respectively.
A power law distribution
$\frac{\partial R}{\partial E}(E)=R_{1}\bigg{(}\frac{E}{E_{1}}\bigg{)}^{-\kappa}$ (31)
where $E_{1}$ is an energy which serves for normalization only yields
$P_{par}=\frac{R_{1}\varepsilon_{l}^{2}}{2-\kappa}\bigg{(}\frac{\epsilon_{l}}{E_{1}}\bigg{)}^{2-\kappa}$ (32)
$S_{p}=2\frac{R_{1}\varepsilon_{l}^{3}}{3-\kappa}\bigg{(}\frac{\epsilon_{l}}{E_{1}}\bigg{)}^{3-\kappa}$ (33)
$\frac{S_{p}}{2P_{par}}=\epsilon_{l}\frac{2-\kappa}{3-\kappa}$ (34)
for $\kappa>3$.
Note that while here we integrate up to infinity, these integrals are essentially unaffected by integrating up to a finite cutoff $\varepsilon_{h}$ as long as $\varepsilon_{h}>>\varepsilon_{l}$ and $\varepsilon_{h}>>\varepsilon$ for exponential distributions. Equivalently, we can say that the ultimate shot noise we measure does not strongly depend on how we cut high energy events, as long as we cut very high energy events which do not belong to the underlying exponential or power law distribution (e.g. saturated events from cosmic rays, high energy LEE). For scale, our $\varepsilon_{l}$ (and $\varepsilon$ in the exponential distribution scenario) should be around the aluminum bandgap $2\Delta_{Al}\approx$ 0.7 meV, and our high energy cutoffs will be on the order of $\sigma_{p}$, i.e. hundreds of meV, decisively putting us in the regime where the severity of our high energy event cut does not strongly set the correlated noise level we measure. We have explicitly tested this by accepting events for the noise analysis with a varying energy cutoff, and have confirmed that the ultimate correlated noise level does not vary with different energy level cuts, as long as the very largest events are cut.
Supplementary Materials Section F Singles Rate over Time
As in Ref. Anthony-Petersen et al. (2025), we have studied the pulse shape of the singles pulses observed in the background dataset and compared them to the aluminum direct hit events in the calibration dataset, and conclude as we did in Ref. Anthony-Petersen et al. (2025) that the background singles events originate within the aluminum films.
In addition to these pulse shape studies, we have measured the rate of singles events over time. To do this, we trigger singles events on a single channel optimal filter, using the averaged pulse shape of above threshold singles events as a template. We then reconstruct the energy these events deposit in our TESs, assuming this same average template, and converting from a current-domain amplitude in the TES as measured by our optimal filter to an energy by applying a current responsivity model $\partial P/\partial I(f)$ as in Refs. Anthony-Petersen et al. (2025); Romani et al. (2024). As described in the main text, we discriminate these events from shared events by comparing the $\chi^{2}$ values for two channel, one amplitude (2x1) optimal filters assuming both singles and shared responses.
Figure 12: (Left) Spectra of singles events observed in the 1 mm and 4 mm detectors. Energies correspond to an absorbed energy in the TES. (Right) Singles rate between 0.7 and 1.8 eV over time, corresponding to the orange band in the left panel.
Comparing the left and right singles spectra for the 1 and 4 mm devices, we see differences in the rate and energy scale of singles observed in different channels. Presumably, some of these differences are due to saturation effects, which would be expected to distort spectra at different energies and relative to each other. As we discuss in Ref. Anthony-Petersen et al. (2025), our singles appear to deposit energy at a single point within a specific QET’s aluminum fins, with energy unevenly shared between the several TESs closest to this event. These different TESs receive different amounts of energy, and therefore saturate at different energies, leading to an artificially broadened and inherently energy dependent pulse shape. We observe evidence of energy dependence in these pulse shapes, and by fitting a pulse model composed of a delta-function and saturated component find that the average singles pulse shape becomes less saturated and more delta-like at lower energies, consistent with this model.
However, some channel-to-channel variation might remain even if the saturation related effects could be compensated for. In principle, some film-to-film variation of singles energy and rate might be expected, given these singles appear to originate with the aluminum QET films. This might be due to, for instance, the different thermal histories of the 1 mm and 4 mm aluminum films, as thicker substrates would be expected to remain cooler during aluminum deposition.
Supplementary Materials Section G Event Rates and Noise over Time
In this letter and supplementary material, we measure the noise and background rate as a function of time in several channels in the 1 mm detector:
  • The correlated noise amplitude.
  • The “uncorrelated” noise rate between 500 Hz and 1 kHz
  • The “uncorrelated” noise rate between 3 and 5 kHz
  • The shared background rate between 2.5 and 5 eV
  • The shared background rate between 5 and 10 eV
  • The singles background rate
To test whether these time varying backgrounds and noise terms may be related, we show fits to the data using simple power law models (see Figs. 13, 14, 15).
Figure 13: Normalized shared and single rates vs. time in the 1 mm detector, with power law fits shown. For shared events between 5 and 10 eV, we show a constant with time model.
Figure 14: Normalized correlated and uncorrelated noise vs. time models in the 1 mm device, with power law fits shown. The color of the dashed fit lines corresponds to the color of the data points shown in the legend.
Figure 15: Normalized correlated noise and shared backgrounds in the 1 mm device between 2.5 and 54 eV, with power law fits shown. These power law fits are statistically inconsistent.
Channel Name Power Law Constant $\kappa$
Correlated Noise $0.635\pm 0.008$
Left Uncorrelated Noise 0.5-1 kHz $0.66\pm 0.07$
Right Uncorrelated Noise 0.5-1 kHz $0.73\pm 0.05$
Left Uncorrelated Noise 3-5 kHz $0.34\pm 0.03$
Right Uncorrelated Noise 3-5 kHz $0.35\pm 0.02$
Shared Backgrounds, 5-10 eV $-0.10\pm 0.12$
Shared Backgrounds, 2.5-5 eV $0.40\pm 0.07$
Left Singles Backgrounds $0.29\pm 0.09$
Right Singles Backgrounds $0.27\pm 0.08$
Table 1: Power law constants, from power law models fit to time series data in different noise and background channels in the 1 mm detector.
Table 1 shows the fit power law constants in the various noise and background channels measured as a function of time. The uncorrelated noises and singles backgrounds have statistically consistent power law constants measured in the left and right channels, consistent with the idea that the two channels observe the same underlying processes. Additionally, we see that below the primary phonon pole the uncorrelated noise is statistically consistent with the correlated noise. This is consistent with the idea, discussed above, that position dependence in the phonon noise spectrum leads to phonon coupled “uncorrelated” noise. However at higher frequencies, we see statistically inconsistent time dependence, suggesting that some other mechanism dominates the noise above several kHz, as discussed above.
In principle, our observed correlated noise could be due to a phonon spectrum that continues to be observable above threshold, however, comparing $\kappa$s for the correlated noise and 2.5-5 eV shared backgrounds shows statistically inconsistent time dependencies. As discussed in the main text, different mechanisms seem required to give time dependence on the meV and eV energy scales.
Finally, we note that the 2.5-5 eV and singles backgrounds appear to have statistically consistent time dependencies. However, our observation of the shared background rate scaling with the detector thickness and the pulse shape evidence for singles originating within the aluminum QET fins strongly suggest that these two event classes have different origins, and that the statistical consistency between their time dependencies is purely coincidental.
Supplementary Materials Section H Noise Comparison to Other Experiments
Here, we describe comparisons of the excess noise observed in the two detectors described in this work to other detectors.
Detector Name Dimensions Holding Correlated Noise Day of Run Noise / Mass Corrected Noise / Mass
(W2/Hz) (W2/(Hz g)) (W2/(Hz g))
This work, 4 mm 1 cm2 $\times$ 4 mm HangingAnthony-Petersen et al. (2024) $3.6\times 10^{-36}$ 12 $3.8\times 10^{-36}$ $3.8\times 10^{-36}$
This work, 1 mm 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) $7.6\times 10^{-37}$ 12 $3.2\times 10^{-36}$ $3.2\times 10^{-36}$
SPICE 1%  [^1] This detector had degraded phonon collection efficiency and therefore lower observed noise due to accidental etching of its W TESs. We correct for this by scaling the noise up by the ratio of the phonon collection efficiencies measured for this work and for the SPICE 1% detector, i.e. (39.1%/28.9%) = 1.33. Anthony-Petersen et al. (2025) 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) $1.0\times 10^{-36}$ 7 $4.4\times 10^{-36}$ $3.1\times 10^{-36}$
CPD v1222Single channel detector, noise reduced by a factor of two for comparison to two channel detectors. Fink et al. (2021); Alkhatib et al. (2021) 7.6 cm Ø$\times$ 1 mm Cirlex clamps $4.2\times 10^{-35}$ Unknown $4.0\times 10^{-36}$ $\sim 2.8\times 10^{-36}$
CPD v2333Not previously published, an updated version of the detector described in Ref. Fink et al. (2021). Single channel detector, noise reduced by a factor of two for comparison to two channel detectors. 7.6 cm Ø$\times$ 1 mm Sapphire clamps 4.7 ×$\times 10^{-35}$ 6 $4.5\times 10^{-36}$ $2.9\times 10^{-36}$
Table 2: Correlated noise amplitude comparison in different low threshold phonon detectors. The two single channel CPD detectors have their noise reduced by a factor of two to correct for the noise being absorbed in a single channel, as opposed to one of two channels. The final column corrects the noise to the expected noise on day 12, assuming the noise falls off with the same power law as measured here ($\kappa=0.63$). For CPDv1, we assume the data was taken on day 7.
Detector Name Dimensions Holding Mass (g) Surface Area (cm2) Sidewall Area (cm2)
This work, 4 mm 1 cm2 $\times$ 4 mm HangingAnthony-Petersen et al. (2024) 0.932 3.6 1.6
This work, 1 mm 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) 0.233 2.4 0.4
SPICE 1% Anthony-Petersen et al. (2025) 1 cm2 $\times$ 1 mm HangingAnthony-Petersen et al. (2024) 0.233 2.4 0.4
CPD v1 Fink et al. (2021); Alkhatib et al. (2021) 7.6 cm Ø$\times$ 1 mm Cirlex clamps 10.6 47.75 2.39
CPD v2444Not previously published, an updated version of the detector described in Ref. Fink et al. (2021). 7.6 cm Ø$\times$ 1 mm Sapphire clamps 10.6 47.75 2.39
CRESST 0.35g Si 2022Angloher et al. (2023) (2 cm)2 $\times$ 0.4 mm Copper clamps 0.35 8.32 0.32
Table 3: Dimensions, weight, and holding of low threshold detectors considered in this work